OpenStudy (anonymous):

I NEED HELP ASAP .. Given the function y = 2/3 x - 10, which set of numbers completes the table? Input..Output -6 -14 -3 0 3 Answer Choices; A.8, 10, 12 B.-13, -12, -11 C.-8, -10, -12 D.-12, -10, -8

OpenStudy (pfenn1):

@iAmDifferent, you can do this. Try plugging in zero for your x and what is y? Come on, you can do this!

OpenStudy (anonymous):

Idk :/

OpenStudy (anonymous):

@pfenn1 , i need this answer like NOW .. My online test is about to time out .

OpenStudy (anonymous):

@ParthKohli , Any help here ?

OpenStudy (pfenn1):

y=2/3 (0)-10=0-10=-10

OpenStudy (anonymous):

So is it A,B,C, or D ?

OpenStudy (pfenn1):

\[y=\frac23(3)-10\]Come on, what is the answer?

OpenStudy (anonymous):

C ?

OpenStudy (pfenn1):

No, so the input of 0 gets you the output of -10, and the input of 3 get you the output of?\[y=\frac23(3)-10=2-10\]

OpenStudy (anonymous):

I got it .. its D :)

OpenStudy (pfenn1):

Correct!!! See! You can do it.

OpenStudy (anonymous):

Thanks :)